OpenStudy (anonymous):

for the simple harmonic motion equation d=9cos((pi/2)t) what is the period

OpenStudy (anonymous):

For cos(t), t going from 0 to 2pi is one period. If I change the function to cos(2t), then t going from 0 to pi makes (2t) go from 0 to 2pi. So now the period is just pi.

OpenStudy (turingtest):

is t in the denominator or the numerator?

OpenStudy (anonymous):

d=9cos((pi/2)t)

OpenStudy (anonymous):

Doesn't it have to be the numerator? Otherwise there would be no discernible period.

OpenStudy (turingtest):

right, but I think your answer is wrong SM... for functions of the form\[A\cos(Bx+C)\]the period is given by \[T=\frac{2\pi}{|B|}\]

OpenStudy (experimentx):

looks like T = 4 ???

OpenStudy (experimentx):

the standard equation of SHM might be \[ y = A\cos(\omega t + \phi)\] And omega might be your angular frequency. you may calculate using analogy between two equations.

OpenStudy (experimentx):

and \[\omega = \frac{2\pi}{T}\] where is T is your time period.

OpenStudy (anonymous):

No, you're right, Turing. I was just giving an example of how it's done. I wasn't working the problem for him/her.

OpenStudy (experimentx):

scooby doovey do .... just give some clue.

OpenStudy (turingtest):

lol^

OpenStudy (kropot72):

\[\omega t=(2\pi f)t=(\pi /2)t\] \[2\pi f=\pi /2\] \[f=(\pi /2)\times(1/2\pi)=(1/4)Hz\] Period=1/f=4seconds

OpenStudy (experimentx):

yes you are right.